Suppose Congress decides to reduce transfer payments (such as welfare) but to increase government purchases of goods and services by an equal amount. That is, it undertakes a change in fiscal policy such that DG 5 −DTR.
a. Would you expect equilibrium income to rise or fall as a result of this change? Why? Check your answer with the following example: Suppose that, initially, c 5 .8, t 5 .25, and Y0 5 600. Now let DG 5 10 and DTR 5 210.
b. Find the change in equilibrium income, DY0.
c. What is the change in the budget surplus, DBS ? Why has BS changed?
1
Expert's answer
2021-08-11T09:18:28-0400


The output of a closed economy is:

"Y=C+I+G+NX"

It can also be written as:

"Y= \bar{C}+c \bar{TR} + \bar{I} + \bar{G} + \bar{N} \bar{X}+ c(1-t)Y"

(a) Total expected equilibrium will increase because of the change as the increase in government purchase will reflect fully in GDP, while consumption will not reduce by equal amount because not all the portion of transfer was used in consumption, some portion was used in savings also.

(b) The change in equilibrium is:

"Δα = \frac{ΔG+cΔTR}{1-c(1-t)} \\

= \frac{10+0.8(-10)}{1-0.8(1-0.25)} \\

= \frac{10-8}{1-0.6} \\

= 5"

(c) The change in budget surplus be:

"ΔBS = tΔY_0 -ΔG -ΔTR \\

= 0.25 \times 5 -10 +10 \\

= 1.25"

The decrease in transfer payment is compensated by increase in government expenditure but still there will be increase in budget surplus because tax revenue has increased due to increased income.